At what temperature would 3.2 g of helium

occupy a volume of 25 L at a pressure of 700

mmHg?


1
Expert's answer
2021-04-01T02:32:51-0400

The ideal gas equation pV = nRT,

where R = 8.31 J/(K·mol)

n = 3.2/4 = 0.8 mole

p = 700 mm Hg = 93,326 Pa

T = 93,326 Pa×0.025 m³/(0.8 mol×8.31 J/(K·mol) ) = 351 K = 78°C

3.2 g of helium occupies a volume of 25 liters at the pressure of 700 mmHg at 78°C.